# Antitrust, Amazon, and Algorithmic Auditing

### v1

@@ Introduction
Next, we shift our focus to product selections facilitated by Amazon’s virtual assistant, Alexa. While many consumers traditionally purchase products from Amazon through its website, there exists an alternative avenue: Alexa enables consumers to make purchases through voice commands, interacting with smart speakers.
@@
Compared to consumers interacting with Amazon on a website, consumers enjoy a significantly smaller action space when interacting with Amazon through a virtual assistant such as Alexa. Rather than presenting a list of search results, Alexa spells out the details of a single product and adds it to the consumer’s cart. By reducing consumer choice, Amazon’s default choices can have a particular strong influence on consumer decisions, even stronger than the selection of a seller in the Buy Box.
Similar to our previous investigations, we start our investigation by collecting information on the products that Alexa selects in response to searches and how these products compare with the search results of a similarly worded search on Amazon’s webpage. Overall, we collected search results for 1,000 different queries from both Amazon Alexa and from Amazon’s website. We also collected the metadata (including product title, ratings, price, and seller information) for the products added to the cart by Alexa and for products appearing among the search results on Amazon’s website. The exact data collection framework is shown in Figure 5 (b). [^6] More details on the data collection can be found in Dash et al. (2022).
We then conducted surveys with 100 participants to elicit consumers’ preferences between the product added to cart by Alexa and the top search result (appearing in the first position) from a desktop search. We selected the top search result for comparison because Amazon identified it as the most relevant product for the query at the time for the customer. We showed our survey participants the titles of the two products, their prices, their average user ratings and the number of ratings received by each of the two products, information we obtained from the Amazon results page during our data collection. We then asked participants the following question: Suppose you are looking for “query string”. Which of the following would you prefer to buy?
@@
This subsection turns to yet another choice architecture by which Amazon may engage in self-preferencing: recommendations on Amazon.
@@
As an e-commerce marketplace, Amazon provides basic product recommendation functions on its platform (see Section 2). It seems reasonable to assume that the results of such recommendations displayed in conjunction with article listings will often be a relevant determinants for consumers’ purchase decisions. These functions may offer Amazon yet another opportunity to “push” its products at the expense of products manufactured and/or sold by others.
Figure 6 shows a screenshot of related item recommendations on Amazon. These related item recommendations appear towards the bottom of almost all Amazon product pages. Naturally, the number of products Amazon can display in this context is limited, and a crucial question is how Amazon’s algorithms decide which products a customer can see.
Amazon displays both (a) organic results, which appear to be determined on the basis of prior customer behavior, and (b) sponsored results, with the items displayed determined through second price auctions in which sellers can participate. [^7] For the example of related item recommendations, see the lower carousel of recommendations in Figure 6. The title of the carousel reads “customer who viewed this item also viewed.” As the title suggests, the items in the carousel are essentially similar to the product on whose product pages they appear based on how customers viewed the items in prior exploration sessions. As has been described elsewhere, these results are curated from item-item collaborative filtering processes (Linden, Smith, and York, 2003; Smith and Linden, 2017)). When Amazon displays sponsored results, it marks those with a tag that reads, on its English language platform, “sponsored.”
@@
To analyze this question, we investigate the products shown on Amazon’s websites as related product recommendations for batteries (Dash et al., 2021). We collected more than 100,000 related item recommendations from the product pages of 5,352 products in the battery category of Amazon using a breadth first search crawler (Dash et al., 2021). Out of these 5,352 products, only 17 products (0.3%) were Amazon private label products.
@@
Another way to look at these numbers is to compare how often private-label products appear as recommended products in the context of organic recommendations vs. sponsored results. Amazon’s own organic recommendation engine recommended the Amazon private-label products from approximately 15% of the product base. Simultaneously, they received recommendations from almost 50% of the product base in the sponsored recommendations.

### v2

@@ Introduction
Next, we shift our focus to product selections facilitated by Amazon’s virtual assistant, Alexa  [^8] Summarized from Dash et al. (2022).. While many consumers traditionally purchase products from Amazon through its website, there exists an alternative avenue: Alexa enables consumers to make purchases through voice commands, interacting with smart speakers.
@@
Compared to consumers interacting with Amazon on a website, consumers enjoy a significantly smaller action space when interacting with Amazon through a virtual assistant such as Alexa. Rather than presenting a list of search results, Alexa spells out the details of a single product and adds it to the consumer’s cart. By reducing consumer choice, Amazon’s default choices can have a particularly strong influence on consumer decisions, even stronger than the selection of a seller in the Buy Box.
Similar to our previous investigations, we start our investigation by collecting information on the products that Alexa selects in response to searches and how these products compare with the search results of a similarly worded search on Amazon’s webpage. Overall, we collected search results for 1,000 different queries from both Amazon Alexa and from Amazon’s website. We also collected the metadata (including product title, ratings, price, and seller information) for the products added to the cart by Alexa and for products appearing among the search results on Amazon’s website. The exact data collection framework is shown in Figure 5 (b).
We then conducted surveys with 100 participants to elicit consumers’ preferences between the product added to cart by Alexa and the top search result (appearing in the first position) from a desktop search. We selected the top search result for comparison because Amazon identified it as the most relevant product for the query at the time for the customer. We showed our survey participants the titles of the two products, their prices, their average user ratings and the number of ratings received by each of the two products information we obtained from the Amazon results page during our data collection. We then asked participants the following question: Suppose you are looking for “query string”. Which of the following would you prefer to buy?
@@
This subsection turns to yet another choice architecture by which Amazon may engage in self-preferencing: recommendations on Amazon  [^9] Summarized from Dash et al. (2021)..
@@
As an e-commerce marketplace, Amazon provides basic product recommendation functions on its platform (see Section 2). It seems reasonable to assume that the results of such recommendations will often be a relevant determinants for consumers’ purchase decisions. These functions may offer Amazon yet another opportunity to “push” its products at the expense of products manufactured and/or sold by others.
Figure 6 shows a screenshot of related item recommendations. These related item recommendations appear towards the bottom of almost all Amazon product pages. Naturally, the number of products Amazon can display in this context is limited, and a crucial question is how Amazon’s algorithms decide which products a customer can see.
Amazon displays both (a) organic results, which appear to be determined on the basis of prior customer behavior, and (b) sponsored results, with the items displayed determined through second-price auctions in which sellers can participate. [^10] For the example of related item recommendations, see the lower carousel of recommendations in Figure 6. The title of the carousel reads “customer who viewed this item also viewed.” As the title suggests, the items in the carousel are essentially similar to the product on whose product pages they appear based on how customers viewed the items in prior exploration sessions. As has been described elsewhere, these results are curated from item-item collaborative filtering processes (Linden, Smith, and York, 2003; Smith and Linden, 2017)). When Amazon displays sponsored results, it marks those with a tag that reads, on its English language platform, “sponsored.”
@@
To analyze this question, we investigate the products shown on Amazon’s websites as related product recommendations for batteries. We collected more than 100,000 related item recommendations from the product pages of 5,352 products in the battery category of Amazon using a breadth first search crawler. Out of these 5,352 products, only 17 products (0.3%) were Amazon private label products.
@@
Another way to look at these numbers is to compare how often private-label products appear as recommended products in the context of organic recommendations vs. sponsored results. Amazon’s own organic recommendation engine recommended Amazon private-label products from approximately 15% of recommendations shown to consumers. Simultaneously, they received recommendations from almost 50% of the product base in the sponsored recommendations.